# Robust DNN Partitioning and Resource Allocation Under Uncertain Inference Time

### v1

@@ VI-B Convergence and Complexity
First, we show the convergence of the proposed algorithms. Fig. 9 illustrates the average number of iterations of Algorithm 1 versus the number of mobile devices. Although the number of iterations of Algorithm 1 cannot be analytically characterized, we can see from Fig. 9 that even when the number of devices $N=30$, Algorithm 1 can terminate after a few iterations. Moreover, the average number of iterations for AlexNet and ResNet152 are similar. In addition, the average number of iterations of Algorithm 1 increases slightly as the number of mobile devices increases significantly. This indicates that Algorithm 1 based on PCCP has better scalability.
@@
Figure 9: The average number of iterations of Algorithm 1 under different numbers of mobile devices in the two models with AlexNet and ResNet152.
@@
Fig. 10(a) and Fig. 10(b) illustrate the convergence trajectories of Algorithm 2 from different initial points in AlexNet and ResNet152 models, respectively. We select three different points as the initial points from all the partitioning points of AlexNet and ResNet152, respectively. For example, the initial points for AlexNet are 3, 7 and 9, while for ResNet152, the initial points are 1, 8 and 9. From Fig. 10, it can be observed that the advantage of using Algorithm 2 is its ability to converge quickly in the early stages of iteration. In addition, Algorithm 2 almost converges to the same objective function value for different initial points.
Figure 10: The convergence trajectories of Algorithm 2 for AlexNet with $D_{n}=220$ ms, and for ResNet152 with $D_{n}=160$ ms.
@@
Then, we show the computational complexity of the proposed algorithms. Fig. 11 illustrates the average runtime of Algorithm 2 on AlexNet and ResNet152. The simulation experiments are implemented using MATLAB and conducted on a laptop computer with an Intel Core i7-8700 3.2 GHz CPU and 16 GB RAM. From Fig. 11, it can be seen that the average runtime of the proposed algorithm increases linearly with the number of mobile devices despite the exponentially growing search space for finding the partitioning decision. Since the ResNet152 model has 10 partitioning points, its average runtime is slightly higher than that of the AlexNet model, which has 9 partitioning points. Combining this with the computational complexity analysis in Section IV, we observed that the complexity of the proposed Algorithm 1 and Algorithm 2 are polynomial time with respect to the number of partitioning points and mobile devices.
@@ VI-C Performance Evaluation
In this subsection, we investigate the impact of different device numbers, risk levels, and task deadlines on the total energy consumption under the AlexNet and ResNet152 models. In addition, we further analyze the violation probability of the task deadline under varying risk levels.
VI-C1 Impact of device numbers
Figure 12: Impact of device numbers for AlexNet with $D_{n}=200$ ms, $B=5$ MHz, and for ResNet152 with $D_{n}=150$ ms, $B=15$ MHz.
@@ VI-C1 Impact of device numbers
Fig. 12 evaluates the impact of the device numbers on total energy consumption. Firstly, we can observe that the total energy consumption increases with the number of mobile devices. Compared to AlexNet, the total energy consumption of ResNet152 rises faster. It is because ResNet152 has a smaller deadline, and to meet the inference time requirement, the mobile devices offload data to the MEC server for execution while simultaneously increasing the local CPU/GPU frequency, which leads to higher energy consumption for both offloading and local computation. Secondly, the performance of the proposed Algorithm 1 is very close to the optimal policy. The computational complexity of the optimal policy is $\mathcal{O}\left(M^{N}\right)$, which is exponential, while the proposed PCCP-based Algorithm 1 can find a stationary point of the DNN model partitioning subproblem, and its computational complexity is polynomial.
VI-C2 Impact of risk levels
@@
Figure 13: The performance of the proposed policy on AlexNet. (a) Energy consumption under different risk levels with $N=12$, $B=10$ MHz and $D_{n}=180$ ms. (b) Energy consumption under different deadlines with $N=12$, $B=10$ MHz and $\varepsilon_{n}=0.02$. (c) Deadline violation probability under different risk levels with $N=12$ and $B=10$ MHz.
@@
Figure 14: The performance of the proposed policy on ResNet152. (a) Energy consumption under different risk levels with $N=12$, $B=30$ MHz and $D_{n}=120$ ms. (b) Energy consumption under different deadlines with $N=12$, $B=30$ MHz and $\varepsilon_{n}=0.04$. (c) Deadline violation probability under different risk levels with $N=12$ and $B=30$ MHz.
@@
Fig. 13(a) and Fig. 14(a) show the total energy consumption at different risk levels. Currently, there is no effective solution that guarantees the deadline for DNN partitioning under inference time uncertainty. For comparison, we consider the worst-case policy using the upper bound of $t_{n,m}^{\mathrm{loc}}$ and $t_{n,m}^{\mathrm{vm}}$ and compare our proposed Algorithm 2 with the worst-case policy.
@@
As shown in Fig. 13(a), the total energy consumption of Algorithm 2 is always lower than that of the worst-case policy. Even when the risk level $\varepsilon=0.02$, total energy consumption can be reduced by nearly $20.7\%$. When the risk level increases to 0.08, the total energy consumption of Algorithm 2 saves $48.3\%$ compared to the worst-case policy. As the risk level increases from 0.02 to 0.08, we observe that the total energy consumption monotonically decreases, which is as expected with (V-B) and (28) that we derived. From (V-B), it can be observed that under a given partitioning decision, $\sigma_{n}$ decreases as $\varepsilon_{n}$ increases, which means that the variance term of the uncertainty in inference time (i.e., the second term on the left side of (V-B)) becomes smaller. Mobile devices can save energy consumption by reducing CPU/GPU frequency. Similarly, it can be seen from (28) that under given communication and computing resources, mobile devices can save energy consumption by selecting appropriate DNN partitioning points.
As shown in Fig. 14(a), ResNet152 exhibits higher energy consumption in Algorithm 2 compared to the worst-case policy when $\varepsilon_{n}$ is small (e.g., 0.02). As discussed in Section IV, the inference time of ResNet152 fluctuates slightly (i.e., its variance is small), while the approximations used in (11) and (12) are conservative. Nevertheless, it is observed that as $\varepsilon_{n}$ increases, the energy consumption of Algorithm 2 gradually becomes lower than that of the worst-case policy. Specifically, Algorithm 2 reduces energy consumption by $2.4\%$ at $\varepsilon_{n}=0.04$ and $8.1\%$ at $\varepsilon_{n}=0.08$.
VI-C3 Impact of task deadlines
@@
Fig. 13(b) and Fig. 14(b) show total energy consumption at various task deadlines for a given risk level. It can be observed that for both AlexNet and ResNet152, the total energy consumption decreases monotonically as task deadlines increase. This is because, as the task deadline increases, mobile devices have more opportunities to select blocks with high computing power requirements and large inference time fluctuations for DNN partitioning and offloading these blocks to the MEC server for execution, thereby reducing local inference energy consumption. In addition, the energy consumption of Algorithm 2 is lower than that of the worst-case policy at different task deadlines. For AlexNet, the energy consumption of Algorithm 2 decreases by $61.7\%$ when the deadline $D_{n}$ varies from 160 ms to 280 ms. Similarly, for ResNet152, the energy consumption of Algorithm 2 decreases by $28.6\%$ when the deadline $D_{n}$ varies from 120 ms to 180 ms.
VI-C4 Deadline violation probability
Leveraging real-world data from Nvidia hardware platforms, we analyze the deadline violation probability of Algorithm 2 under various risk level settings. As illustrated in Fig. 13(c) and Fig. 14(c), we present the deadline violation probabilities at different risk levels. Furthermore, we also provide the violation probabilities for tasks with varying deadlines.
@@ VI-C4 Deadline violation probability
We first observe that the violation probability of Algorithm 2 is always lower than the risk level, which affirms the desired probabilistic guarantees and demonstrates the robustness of Algorithm 2 in handling uncertain DNN inference time. The gap between the risk level and the violation probability can be attributed to the fact that the actual inference time of DNNs does not invariably result in the maximum violation probability. This observation is consistent with our design in Section IV-A, where we approximate the variance of DNN inference time using the maximum value in the CPU/GPU frequency scaling range. Although this approximation introduces some errors, it enhances the robustness of the system. Then, it can be observed that the violation probabilities across different deadlines are quite similar under lower risk levels. However, as the risk level gradually increases, the violation probability associated with larger deadlines tends to be relatively higher, but the violation probability remains smaller than the risk level. The proposed Algorithm 2 can achieve energy savings of nearly $40\%$ and $8\%$ for AlexNet and ResNet152 when the actual violation probability is below $1\%$ (i.e., $\varepsilon_{n}=0.06$).
@@ VII Conclusion
In this paper, we investigated the problem of edge-device collaborative inference under uncertain inference time. Our experiments demonstrate that executing DNN inference tasks on high-performance GPUs can significantly enhance inference speed and reduce variations in inference time. This motivates us to develop an effective scheme for DNN model partitioning and resource allocation to achieve a balance among communication costs, computational requirements, and variations in inference time within edge intelligence systems. Therefore, we formulate the problem as an optimization problem that minimizes the total energy consumption of mobile devices while meeting task probabilistic deadlines. To solve this problem, we employ chance-constrained programming (CCP), which permits occasional violations of the target capacity threshold with a low probability, thereby reformulating the probabilistic constraint problem as a deterministic optimization problem. Then, the optimal solution of local CPU/GPU frequencies and uplink bandwidth allocation and a stationary point of DNN partitioning decisions are obtained using convex optimization and penalty convex-concave procedure (PCCP) techniques, respectively. We evaluate our proposed algorithm with real-world data and widely used DNN models. Extensive simulations demonstrate that the algorithm achieves approximately $40\%$ energy savings for AlexNet and $8\%$ for ResNet152, while maintaining an actual violation probability of less than $1\%$.

### v2

@@ VI-B Convergence and Complexity
First, we show the convergence of the proposed algorithms. Fig. 9 illustrates the average number of iterations of Algorithm 1 versus the number of mobile devices. Although the number of iterations of Algorithm 1 cannot be analytically characterized, we can see from Fig. 9 that even when the number of devices $N=30$, Algorithm 1 can terminate after a few iterations. Moreover, the average number of iterations for ViT-B/32, AlexNet, and ResNet152 are not significantly different. In addition, the average number of iterations of Algorithm 1 increases slightly as the number of mobile devices increases significantly. This indicates that Algorithm 1 based on PCCP has better scalability.
@@
Figure 9: The average number of iterations of Algorithm 1 under different numbers of mobile devices in the three models with ViT-B/32, AlexNet, and ResNet152.
@@
Fig. 10(a) and Fig. 10(b) illustrate the convergence trajectories of Algorithm 2 from different initial points in AlexNet, ResNet152, and ViT-B/32 models, respectively. We select three different points as the initial points from all the partitioning points of AlexNet, ResNet152, and ViT-B/32, respectively. For example, the initial points for AlexNet are 3, 7, and 9; for ResNet152, they are 1, 8, and 9; and for ViT-B/32, they are 3, 4, and 5. From Fig. 10, it can be observed that the advantage of using Algorithm 2 is its ability to converge quickly in the early stages of iteration. In addition, Algorithm 2 almost converges to the same objective function value for different initial points.
Figure 10: The convergence trajectories of Algorithm 2 for AlexNet ($D_{n}=220$ ms), ResNet152 ($D_{n}=160$ ms), and ViT-B/32 ($D_{n}=130$ ms).
@@
Then, we show the computational complexity of the proposed algorithms. Fig. 11 illustrates the runtime of Algorithm 2 on AlexNet, ResNet152, and ViT-B/32. The simulation experiments are implemented using MATLAB and conducted on a laptop computer with an Intel Core i7-8700 3.2 GHz CPU and 16 GB RAM. In Fig. 11, the runtime for AlexNet with $N=6$ is normalized to 1. From Fig. 11, it can be seen that the runtime of the proposed algorithm increases linearly with the number of mobile devices, despite the exponentially growing search space. Since the ResNet152 model has 10 partitioning points, its average runtime is slightly higher than that of the AlexNet model, which has 9 partitioning points. Despite having only 7 partitioning points, the ViT-B/32 model actually exhibits a higher runtime. It is because, as shown in Fig. 10, the ViT-B/32 model requires more iterations than both AlexNet and ResNet152. Combining this with the computational complexity analysis in Section IV, we observed that the complexity of the proposed Algorithm 1 and Algorithm 2 are polynomial time with respect to the number of partitioning points and mobile devices.
In addition, we found that although the theoretical computational complexity of the IPT method is proportional to $M$ and $N$ to the 3.5th power, i.e., $\mathcal{O}\left((NM)^{3.5}\log(1/\xi)\right)$, Fig. 11 shows a much lower practical increase in runtime. Specifically, when the number of devices increases from $N=6$ to $N=36$, the runtimes of AlexNet, ResNet152, and ViT-B/32 increase by less than 5 times. This observed runtime growth is much lower than what would be expected based solely on the theoretical computational complexity. It may be related to factors such as the optimization methods employed by the solver and the performance of the hardware platform. For larger-scale networks, such as $M\gg 10$, and $N\gg 36$, we can consider combining the proposed algorithm with distributed optimization to reduce the computational complexity of the algorithm.
@@ VI-C Performance Evaluation
In this subsection, we investigate the impact of variance approximation, different device numbers, risk levels, and task deadlines on the total energy consumption under the AlexNet, ViT-B/32, and ResNet152 models. In addition, we further analyze the violation probability of the task deadline under varying risk levels.
VI-C1 Impact of variance approximation
The approximation performance of (11) and (12) in Section IV. B is evaluated below. Fig. 12 shows the energy consumption and violation probability of AlexNet, ResNet152, and ViT-B/32 using the maximum and mean variance. The mean variance method is an approximation of (11) and (12). The maximum variance method is approximated as follows:
[…]
where $\forall n\in\mathcal{N},\forall m,m^{\prime}\in\mathcal{M}$.
In Figs. 12 (a), (c), and (e), we first show the violation probability as $\varepsilon_{n}$ increases. It can be observed that the maximum variance method has a lower violation probability than the mean variance method, as the maximum variance method is more conservative. For AlexNet and ResNet152, even when $\varepsilon_{n}$ increases to 0.15, the violation probability of the mean variance method remains lower than the risk level. However, for ViT-B/32, when $\varepsilon_{n}$ exceeds 0.13, the violation probability of the mean variance method is slightly higher than the risk level. Then, in Figs. 12 (b), (d), and (f), we illustrate the changes in energy consumption as $\varepsilon_{n}$ increases. It can be observed that the mean variance method results in lower energy consumption compared to the maximum variance method. For AlexNet, ResNet, and ViT-B/32, the mean variance method can reduce energy consumption by up to $17\%$, $6.8\%$, and $4\%$, respectively, compared to the maximum variance method when $\varepsilon_{n}=0.03$.
(a) Violation probability with AlexNet
(b) Energy consumption with AlexNet
(c) Violation probability with ViT-B/32
(d) Energy consumption with ViT-B/32
(e) Violation probability with ResNet152
(f) Energy consumption with ResNet152
Figure 12: Impact of variance approximation on AlexNet ($N=12$, $B=10$ MHz, $D_{n}=180$ ms), ViT-B/32 ($N=12$, $B=20$ MHz, $D_{n}=100$ ms), and ResNet152 ($N=12$, $B=30$ MHz, $D_{n}=115$ ms).
VI-C2 Performance of Algorithm 1
(a) AlexNet
(b) ResNet152
(c) ViT-B/32
Figure 13: Performance of Algorithm 1 for AlexNet ($D_{n}=200$ ms, $B=5$ MHz), ResNet152 ($D_{n}=150$ ms, $B=15$ MHz), and ViT-B/32 ($D_{n}=130$ ms, $B=10$ MHz).
@@ VI-C2 Performance of Algorithm 1
Fig. 13 evaluates the performance of Algorithm 1 under different DNN models. Firstly, we can observe that the total energy consumption increases with the number of mobile devices. Compared to the random policy, the total energy consumption of Algorithm 1 is much lower. Moreover, as the number of devices increases, the energy consumption of Algorithm 1 increases at a slower rate than the random policy. Secondly, the performance of the proposed Algorithm 1 is very close to the optimal policy. The computational complexity of the optimal policy is $\mathcal{O}\left(M^{N}\right)$, which is exponential, while the proposed PCCP-based Algorithm 1 can find a stationary point of the DNN model partitioning subproblem, and its computational complexity is polynomial.
VI-C3 Impact of risk levels
@@
Figure 14: The performance of the proposed policy on AlexNet ($N=12$ and $B=10$ MHz). (a) Energy consumption under different risk levels with $D_{n}=180$ ms. (b) Energy consumption under different deadlines with $\varepsilon_{n}=0.03$. (c) Deadline violation probability under different risk levels with $D_{n}=180$ ms.
@@
Figure 15: The performance of the proposed policy on ResNet152 ($N=12$ and $B=30$ MHz). (a) Energy consumption under different risk levels with $D_{n}=115$ ms. (b) Energy consumption under different deadlines with $\varepsilon_{n}=0.04$. (c) Deadline violation probability under different risk levels with $D_{n}=115$ ms.
Figure 16: The performance of the proposed policy on ViT-B/32 ($N=12$ and $B=20$ MHz). (a) Energy consumption under different risk levels with $D_{n}=100$ ms. (b) Energy consumption under different deadlines with $\varepsilon_{n}=0.05$. (c) Deadline violation probability under different risk levels with $D_{n}=100$ ms.
@@
Fig. 14(a), Fig. 15(a), and Fig. 16(a) show the total energy consumption at different risk levels. Currently, there is no effective solution that guarantees the deadline for DNN partitioning under inference time uncertainty. For comparison, we compare the worst-case and mean optimization methods with our proposed Algorithm 2.
@@
As shown in Fig. 14(a), the total energy consumption of Algorithm 2 is always lower than that of the worst-case optimization. Even when the risk level $\varepsilon=0.03$, total energy consumption can be reduced by nearly $44.8\%$. When the risk level increases to 0.09, the total energy consumption of Algorithm 2 saves $53.1\%$ compared to the worst-case optimization. As the risk level increases from 0.03 to 0.09, we observe that the total energy consumption monotonically decreases, which is as expected with (V-B) and (28) that we derived. From (V-B), it can be observed that under a given partitioning decision, $\sigma_{n}$ decreases as $\varepsilon_{n}$ increases, which means that the variance term of the uncertainty in inference time (i.e., the second term on the left side of (V-B)) becomes smaller. Mobile devices can save energy consumption by reducing CPU/GPU frequency. Similarly, it can be seen from (28) that under given communication and computing resources, mobile devices can save energy consumption by selecting appropriate DNN partitioning points.
Fig. 15(a) and Fig. 16(a) show that the energy consumption of ResNet152 and ViT-B/32 also gradually decreases as $\varepsilon_{n}$ increases. At $\varepsilon_{n}=0.06$, compared to the worst-case optimization, ResNet152 and ViT-B/32 can save energy consumption by nearly $9.3\%$ and $5.8\%$, respectively. As shown in Figs. 14(c), 15(c), and 16(c), similar to the worst-case optimization, the violation probability of AlexNet and ResNet152 remains close to 0 even when $\varepsilon_{n}=0.06$, while the violation probability of ViT-B/32 is only $3.5\%$. Moreover, we observe that AlexNet on CPU exhibits a higher proportion of energy saving than ResNet152 and ViT-B/32 on GPU. This is because (1) GPUs are more energy-efficient during DNN inference than CPUs, leading to lower relative energy savings when switching to our method, and (2) as discussed in Section IV, the variance in inference time for ResNet152 and ViT-B/32 is relatively small, which limits the potential for our method to further reduce energy consumption. These results demonstrate that our proposed algorithms are particularly effective for DNNs with considerable variations in inference time.
Additionally, it can be seen from Figs. 14(a), 15(a), and 16(a) that the energy consumption of mean-value optimization is lower than that of Algorithm 2, but mean-value optimization lacks probabilistic guarantees. It will be discussed in detail in the following.
VI-C4 Impact of task deadlines
@@
Figs. 14(b), 15(b), and 16(b) show total energy consumption at various task deadlines for a given risk level. It can be observed that for AlexNet, ResNet152, and ViT-B/32, the total energy consumption decreases monotonically as task deadlines increase. This is because, as the task deadline increases, mobile devices have more opportunities to select blocks with high computing power requirements and large inference time fluctuations for DNN partitioning and offloading these blocks to the MEC server for execution, thereby reducing local inference energy consumption. Additionally, the energy consumption of Algorithm 2 is lower than that of the worst-case policy at various task deadlines. For AlexNet, the energy consumption of Algorithm 2 decreases by $54.6\%$ when the deadline $D_{n}$ varies from 160 ms to 280 ms. For ResNet152, the energy consumption of Algorithm 2 decreases by $37.6\%$ when the deadline $D_{n}$ varies from 105 ms to 165 ms. For ViT-B/32, the energy consumption of Algorithm 2 decreases by $17.8\%$ when the deadline $D_{n}$ varies from 90 ms to 150 ms.
VI-C5 Deadline violation probability
Leveraging real-world data from Nvidia hardware platforms, we analyze the deadline violation probability of Algorithm 2 under various risk level settings. As illustrated in Fig. 14(c), 15(c) and 16(c), we present the deadline violation probabilities at different risk levels achieved by Algorithm 2 and the other 2 benchmarks. As expected, the violation probabilities of the worst-case optimization for inference on AlexNet, ResNet152, and ViT-B/32 are all 0. It is because worst-case optimization is the most conservative method, but it also has the highest energy consumption, as shown in Figs. 14(b), 15(b), and 16(b). As for the mean-value optimization, it leads to more than $40.5\%$, $49.5\%$, and $32.7\%$ violation probabilities for AlexNet, ResNet152, and ViT-B/32, respectively. This is due to the fact that the mean-value optimization approach fails to account for uncertainty in inference time. Although it achieves lower energy consumption compared to Algorithm 2, it is a primitive method that bypasses the uncertainty, without any probabilistic guarantees.
@@ VI-C5 Deadline violation probability
Then, we observe that the violation probability of Algorithm 2 is always lower than the risk level, which affirms the desired probabilistic guarantees and demonstrates the robustness of Algorithm 2 in handling uncertain DNN inference time. The gap between the risk level and the violation probability can be attributed to the fact that the actual inference time of DNNs does not invariably result in the maximum violation probability. This observation is consistent with our design in Section IV-A, where we approximate the variance of DNN inference time using the mean value in the CPU/GPU frequency scaling range. Although this approximation introduces some errors, it further reduces energy consumption compared with the maximum variance approximation while showing high robustness. It can be seen that as the risk level gradually increases, although the violation probability rises, it is still far below the risk level. As shown in Figs. 14(a) and 15(a), Algorithm 2 compared with worst-case optimization can achieve nearly $50.4\%$ and $9.4\%$ energy savings for AlexNet and ResNet152, respectively, when the actual violation probability is less than $0.5\%$ (i.e., $\varepsilon_{n}=0.06$). Furthermore, Fig. 16(a) shows that Algorithm 2 achieves about $6.4\%$ energy savings at an actual violation probability of $0.5\%$ (i.e., $\varepsilon_{n}=0.03$).
VI-C6 Performance under dynamic workload
Figure 17: The impact of dynamic workload on inference time and its variation. (a) AlexNet executes local inference without any workload. (b) AlexNet executes local inference under a dynamic workload.
Figure 18: Deadline violation probability and energy consumption under different risk levels in a dynamic workload with $D_{n}=350$ ms, $B=10$ MHz, and $N=12$.
Fig. 17(a) and Fig. 17(b) show the changes in latency and variance of local inference executed by AlexNet before and after the application of a workload. In Fig. 17(a), only AlexNet inference is executed on Jetson Xavier NX CPU, with no other workload. In Fig. 17(b), while the AlexNet inference task is being executed, a trajectory planning task is also added to the CPU for parallel processing. It can be found that the mean inference time and the variation in inference time of AlexNet increase significantly when a workload is added.
Fig. 18 shows the performance of the proposed policy under the dynamic workload. It can be observed that as $\varepsilon_{n}$ increases, energy consumption gradually decreases. Moreover, even when the inference time fluctuates significantly, the actual deadline violation probability is lower than the risk level. This is because, through reasonable model partitioning, the system offloads DNN blocks that are heavily affected by the workload to the MEC server, thereby reducing the impact of high workload on the inference task.
@@ VII Conclusion
In this paper, we investigated the problem of edge-device collaborative inference under uncertain inference time. Our experiments demonstrate that executing DNN inference tasks on high-performance GPUs can significantly enhance inference speed and reduce variations in inference time. This motivates us to develop an effective scheme for DNN model partitioning and resource allocation to achieve a balance among communication costs, computational requirements, and variations in inference time within edge intelligence systems. Therefore, we formulate the problem as an optimization problem that minimizes the total energy consumption of mobile devices while meeting task probabilistic deadlines. To solve this problem, we employ chance-constrained programming (CCP), which permits occasional violations of the target capacity threshold with a low probability, thereby reformulating the probabilistic constraint problem as a deterministic optimization problem. Then, the optimal solution of local CPU/GPU frequencies and uplink bandwidth allocation and a stationary point of DNN partitioning decisions are obtained using convex optimization and penalty convex-concave procedure (PCCP) techniques, respectively. We evaluate our proposed algorithm with real-world data and widely used DNN models. Extensive simulations demonstrate that, compared to worst-case optimization, our proposed algorithm achieves approximately $50.4\%$, $9.4\%$, and $6.4\%$ energy savings for AlexNet, ResNet152, and ViT-B/32, respectively, while maintaining an actual violation probability of not exceeding $0.5\%$.
Finally, we conclude the paper with several promising directions for future research. First, edge intelligence systems involve high-speed mobile scenarios, such as connected vehicles and drones. These mobile devices introduce uncertainties due to their dynamic nature. Consequently, our method can be extended to high-speed mobile scenarios and subject to joint design and optimization. Second, in multi-user edge intelligence systems, edge servers may experience computational resource competition when generating virtual machines (VMs) for each device, thereby introducing additional uncertainty-induced delays. Therefore, joint optimization and allocation of device-edge resources can further enhance system performance. Third, the proposed algorithm can be further extended to cross-continuous inference tasks, such as video streaming. To achieve this, time-related information needs to be incorporated into the system model, and the impact of autocorrelation on inference time uncertainty and the probabilistic constraints in the CCP model should be analyzed.